OpenStudy (anonymous):

given the line, 2x-3y=2 rewrite the equation in 2-point form

OpenStudy (anonymous):

subtract 2x from both sides, in order to keep y on a side by itself; 2x-3y-2x=-2x+2 Simplify to get -3y=-2x+2 Now divide both sides by -3 in order to solve for y (-3y)/-3=(-2x+2)/-3 Simplify once again to get y=(2/3)x-2/3 Now that your function is in the slope intercept form, you can see that m=2/3, and b = -2/3. Therefore; Slope = 2/3 Y-Intercept = (0,-2/3)

OpenStudy (anonymous):

Is it okay @alyannahere

OpenStudy (anonymous):

hi @Rohangrr but how do i do it in 2-point form ? i get 2 other points right? will i just substitute the values ?

ganeshie8 (ganeshie8):

@alyannahere is this the two point form ? lone passing through two points (x1, y1) and (x2, y2) \(y-y1 = \frac{y2-y1}{x2-x1}(x-x1)\)

ganeshie8 (ganeshie8):

*line